The rate expression for the reaction:
5Br-(aq) + BrO3-(aq) + 6H+(aq) → 3Br2(aq) + 3H2O(l)
is Rate = k[Br-] [BrO3-] [H+]2

Assuming that all variables except concentration are kept constant, predict the effect on the rate of reaction if:
a). [Br-] doubles, (other reactant concentrations unchanged)

b). [BrO3-] halves, (other reactant concentrations unchanged)

c). [H+] triples, (other reactant concentrations unchanged)
1
Expert's answer
2020-10-30T06:43:13-0400

a). Rate1 = k[Br-][BrO3-][H+]2

Rate2 = k[2Br-][BrO3-][H+]2

"\frac{Rate1}{Rate2} = \frac{k[Br-][BrO3-][H+]^2}{k[2Br-][BrO3-][H+]^2}"

"\frac{Rate1}{Rate2} = \frac{k[Br-][BrO3-][H+]^2}{2k[Br-][BrO3-][H+]^2} = \frac{1}{2}"

The rate will doubles.

b). Rate1 = k[Br-][BrO3-][H+]2

Rate2 = k[Br-][1/2BrO3-][H+]2

"\frac{Rate1}{Rate2} = \frac{k[Br-][BrO3-][H+]^2}{k[Br-][\frac{1}{2}BrO3-][H+]^2}"

"\frac{Rate1}{Rate2} = \frac{k[Br-][BrO3-][H+]^2}{\frac{1}{2}k[Br-][BrO3-][H+]^2} = 2"

The rate will halves.

c). Rate1 = k[Br-][BrO3-][H+]2

Rate2 = k[Br-][BrO3-][3H+]2

"\frac{Rate1}{Rate2} = \frac{k[Br-][BrO3-][H+]^2}{k[Br-][BrO3-][3H+]^2}"

"\frac{Rate1}{Rate2} = \frac{k[Br-][BrO3-][H+]^2}{9k[Br-][BrO3-][H+]^2} = \frac{1}{9}"

The rate will become 9 times more.